OpenStudy (anonymous):

can you help me solve 7w-8=z solve for w

OpenStudy (saifoo.khan):

w = (8 + z)/7

OpenStudy (anonymous):

7w-8=z add 8 to both sides... 7w=z + 8 divide by w on both sides w = [z + 8]/7